OpenStudy (anonymous):

solve (2x-3) (x+2)=0

OpenStudy (lastdaywork):

The question can be broken into two statements - 2x - 3 = 0 or x + 2 = 0 Can you tell me the answer now ??

OpenStudy (anonymous):

When you have two things multiplied, and their result is equal to zero, it's because one of them was zero. So, there are two possible solutions, either the first factor was zero, or the second one was. In equations: 2x-3=0 x=3/2 or x+2=0 x=-2 The final answer is: x1=3/2 and x2=-2

OpenStudy (y1066380):

(2x-3)(x+2)=0 (2x-3) = 0 OR (x+2) = 0 x=3/2 or x=-2

OpenStudy (lastdaywork):

@fgiancarelli @y1066380 I would like to mention that Open Study follows the policy of "Give Help, Not Answers" Please don't undermine the learning process. Read - http://openstudy.com/code-of-conduct We should give room to the asker to participate in the process of finding solution. Providing complete solution at one go should be avoided.

OpenStudy (skullpatrol):

Well said @LastDayWork

OpenStudy (anonymous):

You are absolutely right, I'm sorry. Won't happen again.